Question #1 Based on the information below, calculate the probability that a child from Joseph and Josephine will have disease K Joseph and Josephine each have a brother and a sister with disease K. Joseph, Josephine and their parents do NOT have the disease. A. 4/9 B. 4/36 C. 2/3 D. 2/36
Not siblings
Answer #1

Q1

Answer: B. 4/36

Since the parents of Josephine and Joseph do not have the disease, they are probably the carriers for the disease K

The probability of Josephine and Joseph being carriers of the disease K is = 2/3

The probability of Josephine and Joseph having a child with disease K is =

=2/3 (Joseph carrier) x 2/3 (Josephine carrier) x ¼ (child that will have K) = 4/36